OpenStudy (anonymous):

(-4/x^2)(sqrt(x^2+64))+(4/x+1) (x(x^2+64)^-1/2))=0 How do you solve for x? I'm having trouble with the algebra!!!

OpenStudy (anonymous):

i know the answer is x=cube root 256. I just don't know how to get there.

OpenStudy (perl):

what do you have so far

OpenStudy (anonymous):

technically this is the furthest i have worked out a minimization problem so now i have to find x to plug it back to the original equation. I don't know how to go further than this

OpenStudy (anonymous):

okay ... multbly x(4/x+1) and you will get (4+x)

OpenStudy (anonymous):

then your eq. (4+x)/(sqr(x^2+64) - 4sqr(x^2+64)/x^2=0

OpenStudy (anonymous):

\[\left( 4+x \right)/\sqrt{x^2+64} - 4\sqrt{X^2 + 64}/ x^2\]

OpenStudy (anonymous):

\[(x^2*(4+x)-4\sqrt{X^2+64}*\sqrt{x^2+64})/(x^2*\sqrt{x^2+64)}\]

OpenStudy (anonymous):

\[4x^2+x^3-4x^2-4*64=0\]

OpenStudy (anonymous):

x^3=4*64=256